The critical angle for total internal reflection for sapphire surrounded by air is 34.4 degree. Calculate the polarizing angle for sapphire.


1
Expert's answer
2021-03-02T18:08:53-0500

Answer

For critical angle for total internal reflection

"\frac{n_s}{n_a}=\frac{sin90^°}{sin\theta_c}"

Where s for sapphire and a for air

"n_s=\frac{1}{sin\theta_c}=\frac{1}{sin34.4^°}"

Now polarization angle

"\theta_p=\tan^{-1}(\frac{n_s}{n_a}) \\=\tan^{-1}(\frac{n_s}{1}) \\=\tan^{-1}(\frac{1}{sin34.4^°}) \\=60.5^°"